### v1

# A Survey of Constraint Formulations in Safe Reinforcement Learning

Abstract
Ensuring safety is critical when applying reinforcement learning (RL) to real-world problems. Consequently, safe RL emerges as a fundamental and powerful paradigm for safely optimizing an agent’s policy from experimental data. A popular safe RL approach is based on a constrained criterion, which solves the problem of maximizing expected cumulative reward under safety constraints. Though there has been recently a surge of such attempts to achieve safety in RL, a systematic understanding of the field is difficult due to 1) the diversity of constraint representations and 2) little discussion of their interrelations. To address this knowledge gap, we provide a comprehensive review of representative constraint formulations, along with a curated selection of algorithms specifically designed for each formulation. Furthermore, we elucidate the theoretical underpinnings that reveal the mathematical mutual relations among common problem formulations. We conclude with a discussion of the current state and future directions of safe reinforcement learning research.
1 Introduction
Reinforcement learning (RL, Sutton and Barto (1998)) is a powerful paradigm for solving sequential decision-making problems through evaluation and improvement, which has made significant progress in many fields such as games (Mnih et al., 2015; Silver et al., 2016), robotics (Levine et al., 2016), and recommendation systems (Afsar et al., 2022). Recently, particular attention has been paid to RL for its use in the fine-tuning of large language models (LLMs) under the name of reinforcement learning from human feedback (RLHF, Ouyang et al. (2022)). RL is a general concept that can be used in many domains and has been spreading to various disciplines as a useful tool.
When applying RL to real problems, safety is an essential requirement (Amodei et al., 2016). Promising areas for safe RL applications include robotics (Pham et al., 2018), autonomous driving (Shalev-Shwartz et al., 2016), and medical care (Jia et al., 2020), and so onL4D. An emerging and crucial use case of safe RL involves refining LLMs using RLHF to align with human preferences. It is critical to prevent harmful outputs (e.g., toxicity, discrimination) while maintaining utility, which is called Safe RLHF (Dai et al., 2023). As such, safe RL has been an active topic in artificial intelligence, which has been studied both theoretically and practically.
Safe RL is a broad concept with different formulations for the different aspects of real-world safety-critical problems. Garcıa and Fernández (2015) provided a survey of safe RL and categorized its optimization criteria into four groups: 1) the constrained criterion (Geibel, 2006), 2) the worst-case criterion (Heger, 1994), 3) the risk-sensitive criterion (Borkar, 2002), and 4) others (e.g., r-squared and value-at-risk). This paper focuses on the constrained criterion of safe RL. It is a growing and powerful group of methods that maximize the expected cumulative reward under safety constraints.
An interesting yet confusing point in the safe RL research based on the constrained criterion is that there are diverse representations of safety constraints yet little discussion regarding the relations among them. Safe RL research in the last decade has focused on developing new algorithms, i.e. how to solve the problem while pursuing performance. New algorithms have continuously been developed under various formulations, thereby making it increasingly challenging to stay abreast of advancements in the field. There are several recent survey papers on safe RL (Brunke et al., 2021; Liu et al., 2021; Gu et al., 2022), but they also focus on methods rather than formulations. As formulation represents the initial phase in comprehending safe RL or implementing algorithms in practical scenarios, it is imperative for the community to comprehensively survey existing literature and lay the groundwork for acquiring a systematic comprehension.
Our contributions.
This paper first provides a comprehensive survey focusing on the constraint formulations in safe reinforcement learning and introduces representative algorithms for each formulation. Also, we discuss the relations between each constraint formulation while defining three notions called transformability, generalizability, and conservative approximation. Specifically, we provide theoretical results that there are two problems called identical or more general safe RL (IoMG-SafeRL) problems, into which other common problems can be transformed. The main contribution of this paper is to bridge the problems faced by the reader with appropriate algorithms by organizing existing research with a focus on constraint formulation.
2 Preliminaries
We consider safe RL problems modeled using a constrained Markov decision process (CMDP, Altman (1999)), which is formally represented as
$\mathcal{M}\cup\mathcal{C}\coloneqq\underbrace{\langle\,\mathcal{S},\mathcal{A},\mathcal{P},H,r,\gamma_{r},\rho\,\rangle}_{\text{Standard MDP}\,(\mathcal{M})}\cup\ \mathcal{C},$ (1)
where $\mathcal{S}$ is a state space $\{s\}$, $\mathcal{A}$ is an action space $\{a\}$, and $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the state transition, where $\mathcal{P}(s^{\prime}\,|\,s,a)$ is the probability of transition from state $s$ to state $s^{\prime}$ when action $a$ is taken. Note that $\Delta(X)$ denotes the probability simplex over the set $X$. In addition, $H\in\mathbb{Z}_{+}$ is the (fixed) finite length of each episode, $r:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the reward function, $\gamma_{r}\in[0,1)$ is the discount factor for the reward, and $\rho\in\Delta(\mathcal{S})$ is the initial state distribution. A key difference from a standard MDP $\mathcal{M}$ lies in the existence of an additional tuple $\mathcal{C}$ to represent safety constraints, which will be referred to as a “constraint tuple” in the rest of this paper. [^1] Though this paper considers finite-horizon discounted CMDPs, key ideas can be extended to infinite and/or undiscounted cases.
A policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ is a function to map from states to distribution over actions. Let $\Pi$ denote a policy class. Given a policy $\pi\in\Pi$, the value function is defined as
$V^{\pi}_{r,h}(s)\coloneqq$ $~{}\mathbb{E}_{\pi}\left[\,\sum_{h^{\prime}=h}^{H}\gamma_{r}^{h^{\prime}}r(s_{h^{\prime}},a_{h^{\prime}})\,\middle|\,s_{h}=s\,\right],$ (2)
where the expectation $\mathbb{E}_{\pi}$ is taken over the random state-action sequence $\{(s_{h^{\prime}},a_{h^{\prime}})\}_{h^{\prime}=h}^{H}$ induced by the policy $\pi$ and the CMDP $\mathcal{M}\cup\mathcal{C}$. Since the initial state $s_{0}$ is sampled from $\rho$, we slightly abuse the notation and define
$V^{\pi}_{r}(\rho)\coloneqq\mathbb{E}_{s\sim\rho}\left[\,V_{r,0}^{\pi}(s)\,\right].$ (3)
Crucially, this paper deals with safe RL involving the “constrained” policy optimization problem. A policy must be within the feasible policy space $\widehat{\Pi}\subseteq\Pi$ that satisfies the safety constraint based on the given constraint tuple $\mathcal{C}$. Therefore, the optimal policy $\pi^{\star}:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ is defined as
$\pi^{\star}\coloneqq\operatorname*{arg\,max}_{\pi\in\widehat{\Pi}}V^{\pi}_{r}(\rho).$ (4)
An overall sequence for solving safe RL problems based on a constrained criterion is illustrated in Figure 1, which consists of 1) problem formulation and 2) policy optimization. An interesting yet complicated point for understanding safe RL research is that there are several formulations for representing the constrained tuple $\mathcal{C}$ and the feasible policy space $\widehat{\Pi}$ depending on the problem settings or applications that researchers have in their minds. In the next section, we will review seven common safety constraint representations that have been well-studied in existing safe RL research.
[…]
Figure 1: A typical sequence for solving safe RL problems based on constrained criteria. Due to the diversity of safety constraint representations and little discussion on their interrelations, it is not easy to understand safe RL research systematically. Unlike existing survey papers that focus on methods, we aim to provide a comprehensive survey from the perspective of formulations on safe RL.
3 Common Constraint Formulations
This paper deals with safe RL based on a constrained criterion. As presented shortly, for a CMDP $\mathcal{M}\cup\mathcal{C}$, common formulations discussed in this paper can be written as
$\max_{\pi\in\Pi}V^{\pi}_{r}(\rho)\quad\text{s.t.}\quad f_{\mathcal{C}}(\pi)\leq 0.$ (5)
In the rest of this paper, we call $f_{\mathcal{C}}:\Pi\rightarrow\mathbb{R}$ a safety constraint function (SCF) and let $\mathcal{F}_{\mathcal{C}}=\{f_{\mathcal{C}}\}$ denote a class of SCFs. Crucially, $\mathcal{C}$ contains the parametric information of the safety constraint; hence, $\mathcal{F}_{\mathcal{C}}$ is regarded as a class of SCFs that can be represented with $\mathcal{C}$. Note that, though $f_{\mathcal{C}}$ often depends on the parameters defined in the standard MDP $\mathcal{M}$ (e.g., $\mathcal{P}$), we omitted it from the notation for simplicity. The feasible policy space $\widehat{\Pi}$ depends on the structure and parameters of $f_{\mathcal{C}}$; hence is, we explicitly represent it as
$\widehat{\Pi}(f_{\mathcal{C}})\coloneqq\{\pi\in\Pi\mid f_{\mathcal{C}}(\pi)\leq 0\}.$ (6)
In the following subsections, we will introduce common safety constraint representations and associated algorithms while organizing based on the aforementioned notations. Representative existing studies and algorithms are summarized according to the problem formulations in Table 1.
3.1 Expected Cumulative Safety Constraint
One of the most popular safe RL formulations is to represent the safety constraint using the same structure of the value function in terms of reward (i.e., $V_{r}^{\pi}(\rho)$). As with the reward function, we first define the value function for safety in a state $s$ at time $h$, denoted as
$V_{c,h}^{\pi}(s)\coloneqq$ $~{}\mathbb{E}_{\pi}\left[\,\sum_{h^{\prime}=h}^{H}\gamma_{c}^{h^{\prime}}c(s_{h^{\prime}},a_{h^{\prime}})\,\middle|\,s_{h}=s\,\right].$
where $c:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is a safety cost function. [^2] Without loss of generality, we assume that the safety cost function $c$ is bounded between 00 and $1$ for all $(s,a)$ as with the reward function $r$. Especially, the assumption that the safety cost is positive is important for our theoretical analyses. We then take expectation with respect to the initial state distribution $\rho$; that is, for a policy $\pi\in\Pi$, we have
$V^{\pi}_{c}(\rho)\coloneqq$ $~{}\mathbb{E}_{s\sim\rho}\left[\,V_{c,0}^{\pi}(s)\,\right].$
Therefore, we define the following safe RL problem based on the value function regarding the safety cost function.
Problem 1.
Let a constraint tuple be $\mathcal{C}\coloneqq\langle\,c,\gamma_{c},\xi\,\rangle$, where $c:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the safety cost function, $\gamma_{c}\in[0,1)$ is the discount factor for the safety, and $\xi\in\mathbb{R}_{+}$ is the safety threshold. Then,
$\max_{\pi}V^{\pi}_{r}(\rho)\quad\text{s.t.}\quad V_{c}^{\pi}(\rho)\leq\xi.$ (7)
Remark 1.
In Problem 1, the SCF $f_{\mathcal{C}}:\Pi\rightarrow\mathbb{R}$ is defined as
$f_{\mathcal{C}}(\pi)\coloneqq V_{c}^{\pi}(\rho)-\xi.$
A reason why Problem 1 is popular is that the safety constraint has a high affinity with the value function in terms of reward. For example, when $\gamma_{r}=\gamma_{c}$, for any $\lambda\in\mathbb{R}$, we have
$V_{r+\lambda c}^{\pi}(\rho)=V_{r}^{\pi}(\rho)+\lambda\cdot V_{c}^{\pi}(\rho),$
which leads to so-called Lagrangian-based methods (Altman, 1999). Thus, since the theoretical properties of the problem have been studied, several algorithms with theoretical guarantees on optimality or safety have been proposed (e.g., Ding et al. (2021), Bura et al. (2022)). Practically, many well-known algorithms, such as constrained policy optimization (CPO, Achiam et al. (2017)) or reward-constrained policy optimization (RCPO, Tessler et al. (2019), are built upon Problem 1.
3.2 State Constraint
Another popular formulation involves leveraging the “state” constraints so that an agent avoids visiting a set of unsafe states, which is well-suited for the case where an autonomous robot needs to behave safely in a hazardous environment (e.g., a disaster site). This type of formulation has been widely adopted by previous studies on safe-critical robotics tasks (Thananjeyan et al., 2021; Thomas et al., 2021; Turchetta et al., 2020; Wang et al., 2023), which is written as follows:
Problem 2.
Let a constraint tuple be $\mathcal{C}\coloneqq\langle\,S_{\text{unsafe}},\gamma_{c},\xi\,\rangle$, where $S_{\text{unsafe}}\subseteq\mathcal{S}$ is the set of unsafe states, $\gamma_{c}\in[0,1)$ is the discount factor for the safety cost function, and $\xi\in\mathbb{R}_{+}$ is the safety threshold. Then,
$\max_{\pi}V^{\pi}_{r}(\rho)\quad\text{s.t.}\quad\mathbb{E}_{\pi}\left[\,\sum_{h=0}^{H}\gamma_{c}^{h}\mathbb{I}(s_{h}\in S_{\text{unsafe}})\,\right]\leq\xi,$
where $\mathbb{I}(\cdot)$ is the indicator function.
Remark 2.
In Problem 2, the SCF $f_{\mathcal{C}}:\Pi\rightarrow\mathbb{R}$ is defined as
$f_{\mathcal{C}}(\pi)\coloneqq\mathbb{E}_{\pi}\left[\,\sum_{h=0}^{H}\gamma_{c}^{h}\mathbb{I}(s_{h}\in S_{\text{unsafe}})\,\right]-\xi.$
3.3 Joint Chance Constraint
Policy optimization under joint chance constraints has been studied especially in the field of control theory such as Ono et al. (2015) and Pfrommer et al. (2022). This type of safety-constrained policy optimization problem is typically formulated as follows:
Problem 3.
Let a constraint tuple be $\mathcal{C}\coloneqq\langle\,S_{\text{unsafe}},\xi\,\rangle$, where $S_{\text{unsafe}}\subseteq\mathcal{S}$ is the set of unsafe states and $\xi\in\mathbb{R}_{+}$ is the safety threshold. Then,
$\max_{\pi}V^{\pi}_{r}(\rho)\quad\text{s.t.}\quad\mathbb{P}_{\pi}\left[\,\bigvee_{h=0}^{H}s_{h}\in S_{\text{unsafe}}\,\right]\leq\xi,$
where $\mathbb{P}_{\pi}$ represents a probability that is computed over the random state-action sequences $\{(s_{h^{\prime}},a_{h^{\prime}})\}_{h^{\prime}=h}^{H}$ induced by the policy $\pi$ and the CMDP $\mathcal{M}\cup\mathcal{C}$.
Remark 3.
In Problem 3, the SCF $f_{\mathcal{C}}:\Pi\rightarrow\mathbb{R}$ is defined as
$f_{\mathcal{C}}(\pi)\coloneqq\mathbb{P}_{\pi}\left[\,\bigvee_{h=0}^{H}s_{h}\in S_{\text{unsafe}}\,\right]-\xi.$
It is quite challenging to directly solve the Problem 3 characterized by joint chance-constraints; thus, most of the previous work does not directly deal with this type of constraint and uses some approximations or assumptions. For example, Pfrommer et al. (2022) assume a known linear time-invariant dynamics. Also, Ono et al. (2015) conservatively approximate the joint chance constraint into the constraint with an additive structure as in Problem 2 using the following inequality (for more details, see the proof of Theorem 2):
$\mathbb{P}_{\pi}\left[\,\bigvee_{h=1}^{H}s_{h}\in S_{\text{unsafe}}\,\right]\leq\mathbb{E}_{\pi}\left[\,\sum_{h=1}^{H}\mathbb{I}(s_{h}\in S_{\text{unsafe}})\,\right].$
3.4 Expected Instantaneous Safety Constraint with Time-variant Threshold
Though Problems 1, 2, and 3 formulate the long-term safety constraint with additive structures or joint chance constraints, there are a few papers (e.g., Pham et al. (2018)) that focus on “instantaneous” safety constraint. In this problem, an agent is required to satisfy a safety constraint at every time step.
Problem 4.
Let a constraint tuple be $\mathcal{C}\coloneqq\langle\,c,\{\xi_{h}\}_{h=0}^{H}\,\rangle$, where $c:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the safety cost function, and $\xi_{h}\in\mathbb{R}_{+}$ is the safety threshold for time step $h\in[H]$. Then,
$\max_{\pi}V^{\pi}_{r}(\rho)\quad\text{s.t.}\quad\mathbb{E}_{\pi}\left[\,c(s_{h},a_{h})\,\right]\leq\xi_{h},\ \forall h\in[H].$
Remark 4.
In Problem 4, the SCF $f_{\mathcal{C}}:\Pi\rightarrow\mathbb{R}$ is defined as
$f_{\mathcal{C}}(\pi)\coloneqq\max_{h\in[H]}\Bigl{(}\mathbb{E}_{\pi}\left[\,c(s_{h},a_{h})\,\right]-\xi_{h}\Bigr{)}.$
3.5 Almost Surely Cumulative Safety Constraint
Unlike Problem 1 where the expectation $\mathbb{E}_{\pi}$ regarding the safety constraint is taken, we often want to guarantee safety almost surely (i.e., probability of $1$). This problem has been recently studied in Sootla et al. (2022a) and Sootla et al. (2022b), which is based on a stricter safety notion. In such cases, the safe RL problem is formulated as follows:
Problem 5.
Let a constraint tuple be $\mathcal{C}\coloneqq\langle\,c,\gamma_{c},\xi\,\rangle$, where $c:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the safety cost function, $\gamma_{c}\in[0,1)$ is the discount factor for the safety, and $\xi\in\mathbb{R}_{+}$ is the safety threshold. Then,
$\max_{\pi}V^{\pi}_{r}(\rho)\quad\text{s.t.}\quad\mathbb{P}_{\pi}\left[\,\sum_{h=0}^{H}\gamma_{c}^{h}c(s_{h},a_{h})\leq\xi\,\right]=1.$
Remark 5.
In Problem 5, the SCF $f_{\mathcal{C}}:\Pi\rightarrow\mathbb{R}$ is defined as
$f_{\mathcal{C}}(\pi)\coloneqq 1-\mathbb{P}_{\pi}\left[\,\sum_{h=0}^{H}\gamma_{c}^{h}c(s_{h},a_{h})\leq\xi\,\right].$
3.6 Almost Surely Instantaneous Safety Constraint with Time-invariant Threshold
Some existing studies formulate safe RL problems via an instantaneous constraint, attempting to ensure safety even during the learning stage while aiming for extremely safety-critical applications such as planetary exploration (Turchetta et al., 2016; Wachi et al., 2018; Wachi and Sui, 2020; Wang et al., 2023). Such studies require the agent to satisfy the following instantaneous safety constraint at every time step.
Problem 6.
Let a constraint tuple be $\mathcal{C}\coloneqq\langle\,c,\xi\,\rangle$, where $c:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the safety cost function, and $\xi\in\mathbb{R}_{+}$ is the threshold. Then,
$\max_{\pi}V^{\pi}_{r}(\rho)\quad\text{s.t.}\quad\mathbb{P}_{\pi}\left[\,c(s_{h},a_{h})\leq\xi\,\right]=1,\ \forall h\in[H].$
Remark 6.
In Problem 6, the SCF $f_{\mathcal{C}}:\Pi\rightarrow\mathbb{R}$ is defined as
$f_{\mathcal{C}}(\pi)\coloneqq 1-\prod_{h=0}^{H}\mathbb{P}_{\pi}\left[\,c(s_{h},a_{h})\leq\xi\,\right].$
This formulation is also related to notions in the control theory, called control barrier functions (Ames et al., 2016; Cheng et al., 2019) or Lyapunov functions (Berkenkamp et al., 2017). Note that while these functions are typically required to be positive, it is possible to use the same formulation in Problem 6 by defining them as $-c$ and setting $\xi=0$.
3.7 Almost Surely Instantaneous Safety Constraint with Time-variant Threshold
As a similar formulation to Problem 6, Wachi et al. (2023). has recently introduced a problem called the generalized safe exploration (GSE) problem, which is written as follows:
Problem 7.
Let a constraint tuple be $\mathcal{C}\coloneqq\langle\,c,\{\xi_{h}\}_{h=0}^{H}\,\rangle$, where $c:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the safety cost function, and $\xi_{h}\in\mathbb{R}_{+}$ is the safety threshold for time step $h\in[H]$. Then,
$\max_{\pi}V^{\pi}_{r}(\rho)\quad\text{s.t.}\quad\mathbb{P}_{\pi}\left[\,c(s_{h},a_{h})\leq\xi_{h}\,\right]=1,\forall h\in[H].$
Remark 7.
In Problem 7, the SCF $f_{\mathcal{C}}:\Pi\rightarrow\mathbb{R}$ is defined as
$f_{\mathcal{C}}(\pi)\coloneqq 1-\prod_{h=0}^{H}\mathbb{P}_{\pi}\left[\,c(s_{h},a_{h})\leq\xi_{h}\,\right].$
This formulation is quite similar to Problem 6 with the only difference being that the safety threshold $\xi_{h}$ is time-variant. An apparent benefit of Problem 7 compared to Problem 6 is that we can cover a wider range of applications such that the speed limit changes during driving. Additionally, it goes beyond that and offers us more important advantages regarding the theoretical relations with Problem 5, which we will present shortly in Theorem 3.
3.8 Other Constrained Formulations
In the line of safe RL work based on constraints, various attempts were left unexplored and not integrated into our theoretical framework within this paper.
A notable instance of such problem formulations defines a safety constraint using the variance of the return (Tamar et al., 2012) which is represented as
$\max_{\pi}V^{\pi}_{r}(\rho)\quad\text{s.t.}\quad\textrm{Var}\left[\,V_{r}^{\pi}(\rho)\,\right]<\xi.$
Conventional RL algorithms depend on ergodicity assumption; that is, any state $s$ is eventually reachable from any other state $s^{\prime}$ by following a suitable policy. This assumption does not hold in many practical applications since an agent cannot recover on its own after catastrophic actions. Moldovan and Abbeel (2012) removed the ergodicity assumption and proposed an algorithm in which an agent is required to guarantee returnability to the initial safe state.
While this paper has dealt with numerical safety constraints, there have been attempts to represent them using a certain language. For example, Fulton and Platzer (2018) represents the safety constraint via formal languages. This constraint representation is quite useful for leveraging human knowledge in RL, which leads to powerful solutions such as shielding methods (Alshiekh et al., 2018). Finally, Yang et al. (2021) uses natural language to represent safety constraints. This approach is friendly for the general public, which would be a promising method for applying safe RL to real problems.
Problem Type Representative Work Algorithm Theoretical Guarantee Open Source Software (OSS)
Optimality Safety
Problem 1 Online Achiam et al. (2017) CPO -- -- A, SSA, FSRL, SafePO, OmniSafe
Ray et al. (2019) TRPO-Lagrangian -- -- A, SSA, FSRL, SafePO, OmniSafe
PPO-Lagrangian -- -- A, SSA, FSRL, SafePO, OmniSafe
Tessler et al. (2019) RCPO -- -- A, SafePO, OmniSafe
Liu et al. (2020) IPO -- -- A, OmniSafe
Yang et al. (2020) PCPO -- -- A, SafePO, OmniSafe
Stooke et al. (2020) PID-Lagrangian -- -- A, SafePO, OmniSafe
Zhang et al. (2020) FOCOPS -- -- A, FSRL, SafePO, OmniSafe
Ding et al. (2020) NPG-PD Y C --
Bharadhwaj et al. (2021) CSC -- -- A
Ding et al. (2021) OPDOP Y C --
Bai et al. (2022) CSPDA Y C --
As et al. (2021) LAMBDA -- -- A
Xu et al. (2021) CRPO Y C OmniSafe
Yu et al. (2022) SEditor -- -- A
Bura et al. (2022) DOPE Y T and C --
Liu et al. (2022) CVPO Y C A, FSRL
Zhang et al. (2022) P3O -- -- A, OmniSafe
Offline Le et al. (2019) CBPL -- T and C A
Lee et al. (2021) COptiDICE -- T A, OSRL, OmniSafe
Wu et al. (2021) CMOMDPs Y T and C --
Xu et al. (2022) CPQ -- T A, OSRL
Liu et al. (2023b) CDT -- T A, OSRL
Problem 2 Online Turchetta et al. (2020) CISR -- -- A
Thomas et al. (2021) SMBPO -- C A
Thananjeyan et al. (2021) Recovery RL -- -- A
Wang et al. (2023) -- -- T and C A
Problem 3 Online Ono et al. (2015) CCDP -- T and C --
Pfrommer et al. (2022) -- Y T and C --
Mowbray et al. (2022) -- -- T and C A
Kordabad et al. (2022) -- -- T and C --
Problem 4 Online Pham et al. (2018) OptLayer -- T and C A
Amani et al. (2021) SLUCB Y T and C --
Offline Amani and Yang (2022) Safe-DPVI Y T and C --
Problem 5 Online Sootla et al. (2022b) Sauté RL Y C A, SafePO, OmniSafe
Sootla et al. (2022a) Simmer RL Y C A, SafePO, OmniSafe
Problem 6 Online Turchetta et al. (2016) SafeMDP -- T and C A
Berkenkamp et al. (2017) SMbRL -- T and C A
Fisac et al. (2018) -- -- T and C --
Wachi et al. (2018) SafeExpOpt-MDP -- T and C A
Dalal et al. (2018) SafeLayer -- T and C A
Cheng et al. (2019) RL-CBF -- T and C A
Wachi and Sui (2020) SNO-MDP Y T and C A
Wang et al. (2023) -- -- C --
Problem 7 Online Shi et al. (2023) LSVI-NEW Y T and C --
Wachi et al. (2023) MASE Y T and C A
Table 1: Common safe RL formulations based on the constrained criterion and associated representative work. Type indicates whether each safety RL is based on online or offline RL settings. In the Theoretical Guarantee column, Y indicates the (near-)optimality of the policy obtained by an algorithm. Also, T means that safety is guaranteed during training, and C means that safety is guaranteed after convergence. Note that offline algorithms are inherently safe during training since there is no interaction between the agent and the environment. In the OSS column, A means a public authors’ implementation exists, and SSA is an abbreviation of the Safety Starter Agent repository (Ray et al. (2019), https://github.com/openai/safety-starter-agents). Also, FSRL (Liu et al. (2023a), https://github.com/liuzuxin/FSRL), OSRL (Liu et al. (2023a), https://github.com/liuzuxin/OSRL), SafePO (Ji et al. (2023), https://github.com/PKU-Alignment/Safe-Policy-Optimization), and OmniSafe (Ji et al. (2023), https://github.com/PKU-Alignment/omnisafe) are recent and actively maintained repositories for online and offline safe RL, which will lead to the ease of the process of adopting safe RL algorithms.
Problem 1 ($\mathbb{E}_{\pi}$ + Cumulative) $\bullet\ \mathcal{C}=\langle\,c,\gamma_{c},\xi\,\rangle$ $\bullet\ f_{\mathcal{C}}(\pi)=V_{c}^{\pi}(\rho)-\xi$ Problem 2 ($\mathbb{E}_{\pi}$ + Cumulative) $\bullet\ \mathcal{C}=\langle\,S_{\text{unsafe}},\gamma_{c},\xi\,\rangle$ $\bullet\ f_{\mathcal{C}}(\pi)=\mathbb{E}_{\pi}\left[\,\sum_{h=0}^{H}\gamma_{c}^{h}\mathbb{I}(s_{h}\in S_{\text{unsafe}})\,\right]-\xi$ Problem 3 (Joint Chance Constraint) $\bullet\ \mathcal{C}=\langle\,S_{\text{unsafe}},\xi\,\rangle$ $\bullet\ f_{\mathcal{C}}(\pi)\coloneqq\mathbb{P}_{\pi}\left[\,\bigvee_{h=0}^{H}s_{h}\in S_{\text{unsafe}}\,\right]-\xi$ Problem 4 ($\mathbb{E}_{\pi}$ + Instantaneous) $\bullet\ \mathcal{C}=\langle\,c,\{\xi_{h}\}_{h=0}^{H}\,\rangle$ $\bullet\ f_{\mathcal{C}}(\pi)\coloneqq\max_{h\in[H]}\Bigl{(}\mathbb{E}_{\pi}\left[\,c(s_{h},a_{h})\,\right]-\xi_{h}\Bigr{)}$ TransformableConservative ApproximationTransformableTransformableConservative Approximation
Figure 2: Relations among common safe RL formulations based on $\mathbb{E}_{\pi}$ and the one with chance constraints.
Problem 5 ($\mathbb{P}_{\pi}$ + Cumulative) $\bullet\ \mathcal{C}=\langle\,c,\gamma_{c},\xi\,\rangle$ $\bullet\ f_{\mathcal{C}}(\pi)=1-\mathbb{P}_{\pi}\left[\,\sum_{h=0}^{H}\gamma_{c}^{h}c(s_{h},a_{h})\leq\xi\,\right]$ Problem 6 ($\mathbb{P}_{\pi}$ + Instantaneous) $\bullet\ \mathcal{C}=\langle\,c,\xi\,\rangle$ $\bullet\ f_{\mathcal{C}}(\pi)=1-\prod_{h=0}^{H}\mathbb{P}_{\pi}\left[\,c(s_{h},a_{h})\leq\xi\,\right]$ Problem 7 ($\mathbb{P}_{\pi}$ + Instantaneous) $\bullet\ \mathcal{C}=\langle\,c,\{\xi_{h}\}_{h=0}^{H}\,\rangle$ $\bullet\ f_{\mathcal{C}}(\pi)\coloneqq 1-\prod_{h=0}^{H}\mathbb{P}_{\pi}\left[\,c(s_{h},a_{h})\leq\xi_{h}\,\right]$ TransformableTransformable
Figure 3: Relations among common safe RL formulations based on $\mathbb{P}_{\pi}$ (i.e., almost-surely constraints).
4 Theoretical Relations Among Common Constraint Formulations of Safe RL
In this section, we provide theoretical understandings regarding the interrelations among common safe RL formulations presented in Section 3.
4.1 Definitions
We first introduce and define two important notions called transformability and generalizability. Also, we define a notion called conservative approximation.
Definition 1 (Transformability).
Let $\mathcal{M}\cup\mathcal{C}_{1}$ and $\mathcal{M}\cup\mathcal{C}_{2}$ denote two CMDPs that are respectively characterized by different constraint tuples $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. Let $\mathcal{F}_{\mathcal{C}_{1}}$ and $\mathcal{F}_{\mathcal{C}_{2}}$ denote two different classes of SCFs based on $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, respectively. For any SCF $f_{\mathcal{C}_{1}}\in\mathcal{F}_{\mathcal{C}_{1}}$, if there exists $f_{\mathcal{C}_{2}}\in\mathcal{F}_{\mathcal{C}_{2}}$ such that
$\widehat{\Pi}(f_{\mathcal{C}_{1}})=\widehat{\Pi}(f_{\mathcal{C}_{2}}),$
then we say that the problem characterized by $\mathcal{M}\cup\mathcal{C}_{1}$ can be transformed into that characterized by $\mathcal{M}\cup\mathcal{C}_{2}$.
Definition 2 (Generalizability).
Let $N\in\mathbb{Z}_{+}$ denote a positive integer. Let $\{\mathcal{M}\cup\mathcal{C}_{i}\}_{i=1}^{N}$ denote a set of $N$ CMDPs that are respectively characterized by different constraint tuples $\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{N}$. Suppose that, for all $i\in[N]$, the problem characterized by $\mathcal{M}\cup\mathcal{C}_{i}$ can be transformed into that by $\mathcal{M}\cup\widetilde{\mathcal{C}}$. Then, we call the problem characterized by $\mathcal{M}\cup\widetilde{\mathcal{C}}$ is an identical or more general safe RL (IoMG-SafeRL) problem over the set of problems respectively characterized by $\mathcal{M}\cup\mathcal{C}_{1},\mathcal{M}\cup\mathcal{C}_{2},\ldots,\mathcal{M}\cup\mathcal{C}_{N}$.
Definition 3 (Conservative Approximation).
Let $\mathcal{M}\cup\mathcal{C}_{1}$ and $\mathcal{M}\cup\mathcal{C}_{2}$ be two CMDPs that are respectively characterized by different constraint tuples $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. Let $\mathcal{F}_{\mathcal{C}_{1}}$ and $\mathcal{F}_{\mathcal{C}_{2}}$ respectively denote two classes of SCFs based on $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. For any SCF $f_{\mathcal{C}_{1}}\in\mathcal{F}_{\mathcal{C}_{1}}$, if there exists $f_{\mathcal{C}_{2}}\in\mathcal{F}_{\mathcal{C}_{2}}$ such that
$\widehat{\Pi}(f_{\mathcal{C}_{1}})\supset\widehat{\Pi}(f_{\mathcal{C}_{2}}),$
then we say that the problem characterized by $\mathcal{M}\cup\mathcal{C}_{2}$ is a conservative approximation of that characterized by $\mathcal{M}\cup\mathcal{C}_{1}$.
4.2 Preliminary Lemmas
To present the main theoretical results, we first list three necessary lemmas as preliminaries. The first lemma is based on Sootla et al. (2022b) and Wachi et al. (2023), which describes a theoretical connection between additive and instantaneous safety constraints.
Lemma 1.
Define a new variable $\eta_{h}$ meaning the remaining safety budget associated with the discount factor $\gamma_{c}$ such that
$\eta_{h}\coloneqq\gamma_{c}^{-h}\cdot\left(\xi-\sum_{h^{\prime}=0}^{h-1}\gamma_{c}^{h^{\prime}}c(s_{h^{\prime}},a_{h^{\prime}})\right),\ \forall h\in[H].$ (8)
Then, the following relation between additive and instantaneous constraints holds:
$\sum_{h^{\prime}=0}^{h-1}\gamma_{c}^{h^{\prime}}c(s_{h^{\prime}},a_{h^{\prime}})\leq\xi\iff c(s_{h},a_{h})\leq\eta_{h},\quad\forall h\in[H].$
Proof.
By definition, the new variable $\eta_{h}$ satisfies the following recurrence formula:
$\eta_{h+1}=\gamma_{c}^{-1}\cdot\left(\,\eta_{h}-c(s_{h},a_{h})\,\right)\quad\text{with}\quad\eta_{0}=\xi.$ (9)
Combining (9) and $c(s,a)\geq 0,\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, we have
$\eta_{H+1}\geq 0\iff\eta_{h}-c(s_{h},a_{h})\geq 0,\quad\forall h\in[H].$
By definition of $\eta_{h}$, we have
$\sum_{h^{\prime}=0}^{H}\gamma_{c}^{h^{\prime}}c(s_{h^{\prime}},a_{h^{\prime}})\leq\xi\iff\eta_{H+1}\geq 0.$ (10)
Therefore, we obtain the desired lemma. ∎
Lemma 2.
Problem 1 can be transformed into Problem 4.
Proof.
By applying Lemma 1, the safety constraint of Problem 1 satisfies
$V_{c}^{\pi}(\rho)\leq\xi\iff\mathbb{E}_{\pi}\left[\,c(s_{h},a_{h})\,\right]\leq\mathbb{E}_{\pi}\left[\,\eta_{h}\,\right],\ h\in[H].$
Therefore, Problem 1 is identical to a special case of Problem 4 with $\xi_{h}\coloneqq\mathbb{E}_{\pi}[\eta_{h}]$ for all $h\in[H]$. Hence, we obtain the desired lemma. ∎
Lemma 3.
Problem 5 can be transformed into Problem 7.
Proof.
By applying Lemma 1, we have
$\mathbb{P}_{\pi}\left[\,\sum_{h=0}^{H}\gamma_{c}^{h}c(s_{h},a_{h})\leq\xi\,\right]=1$
$\iff\mathbb{P}_{\pi}\left[\,c(s_{h},a_{h})\leq\eta_{h}\,\right]=1,\quad\forall h\in[H].$
Therefore, Problem 5 is identical to a special case of Problem 7 with $\xi_{h}\coloneqq\eta_{h}$ for all $h\in[H]$. Therefore, we obtain the desired lemma. ∎
4.3 Two IoMG-SafeRL Problems
We now provide three theorems on the interrelations among the common safe RL problems. Crucially, we show that Problems 4 and 7 can be regarded as two IoMG-SafeRL problems of other problems. We also show that Problem 4 with $\gamma_{c}=1$ is a conservative approximation of Problem 3. Conceptual illustrations are given in Figures 2 and 3.
Theorem 1.
Problem 4 is an IoMG-SafeRL problem over Problems 1 and 2.
Proof.
By Lemma 2, Problem 1 can be transformed into Problem 4. Also, Problem 2 can be easily transformed into Problem 1 by defining $c(s,a)\coloneqq\mathbb{I}(s\in S_{\text{unsafe}})$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. In summary, Problem 4 is an IoMG-SafeRL problem over Problems 1 and 2. ∎
Theorem 2.
Problem 4 with $\gamma_{c}=1$ is a conservative approximation of Problem 3.
Proof.
This lemma mostly follows from Theorem 1 in Ono et al. (2015). Regarding the constraint in Problem 3, we have the following chain of equations:
$\mathbb{P}_{\pi}\left[\,\bigvee_{h=1}^{H}s_{h}\in S_{\text{unsafe}}\,\right]$ $\ \leq\sum_{h=1}^{H}\mathbb{P}_{\pi}\left[\,s_{h}\in S_{\text{unsafe}}\,\right]$
$\ =\sum_{h=1}^{H}\mathbb{E}_{\pi}\left[\,\mathbb{I}(s_{h}\in S_{\text{unsafe}})\,\right]$
$\ =\mathbb{E}_{\pi}\left[\,\sum_{h=1}^{H}\mathbb{I}(s_{h}\in S_{\text{unsafe}})\,\right].$
In the first step, we used Boole’s inequality (i.e., $\Pr[A\cup B]\leq\Pr[A]+\Pr[B]$). The final term is the left-hand side of the constraint in Problem 2 with $\gamma_{c}=1$, which implies that Problem 2 is a conservative approximation of Problem 3. Therefore, we obtained the desired theorem. ∎
Theorem 3.
Problem 7 is an IoMG-SafeRL problem over Problems 5 and 6.
Proof.
By Lemma 3, Problem 5 can be transformed into Problem 7. Also, Problem 6 is a special case of Problem 7 where $\xi_{h}$ is a constant for all $h\in[H]$. Hence, Problem 7 is an IoMG-SafeRL problem over Problems 5 and 6. ∎
5 Discussion
We conclude by discussing the current state and future directions of safe RL based on constrained criteria.
5.1 Formulation and Algorithm Selection
As we have discussed in this paper, when we adopt the safe RL paradigm, there are two main steps: 1) problem formulation and 2) policy optimization. Based on the survey results, let us discuss how to solve safe RL problems via constraints.
Problem formulation.
As presented in Section 3, constraint formulations in safe RL are divided into two classes: one based on $\mathbb{E}_{\pi}$ and the other based on $\mathbb{P}_{\pi}$. Constraint formulations with expectation $\mathbb{E}_{\pi}$ (i.e., Problems 1 - 4) represent the safety constraints using the expected value, which focuses on the “averaged” performance of safety. On the other hand, constraint formulations based on $\mathbb{P}_{\pi}$ (i.e., Problems 5 - 7) require an agent to guarantee safety almost surely. When the problem is formulated based on $\mathbb{P}_{\pi}$, the achieved safety level is usually higher by nature. Still, there is also a drawback that the reward performance is usually lower due to stricter safety constraints. Which formulation to adopt should depend on which level of safety is faced with the problem.
Policy optimization.
When optimizing a policy, we must select a proper algorithm that corresponds to the problem formulation. Given the current situation, the easiest way is to use algorithms implemented in well-used OSS such as FSRL/OSRL (Liu et al., 2023a) and SafePO/OmniSafe (Ji et al., 2023). Note that, as shown in Table 1, the algorithms implemented in the above OSS are mostly based on Problems 1 and 5. In the long run, however, the algorithms based on Problems 4 and 7 (i.e., IoMG-SafeRL problems) are also promising since instantaneous constraints are easier to handle than cumulative ones both theoretically and empirically.
When should your agent be safe?
Another perspective in the algorithm selection should include when an agent needs to satisfy the safety constraint. Existing algorithms can be divided into two classes (see “safety” column in Table 1). The first class (w/o T) tries to achieve the required level of safety after convergence while encouraging safety during training. The algorithms in this class are usually based on Problems 1 or 5 where the safety constraints are represented using the additive safety cost structure. The second class (w/ T) tries to guarantee safety even during training as well as after convergence. The algorithms in this class are typically built upon Problems 6 or 7 where the safety constraints are instantaneous. Which class of algorithm you should choose depends solely on the problem and application to be addressed.
5.2 Online RL vs. Offline RL
As shown in Table 1, most of the existing safe RL literature considers “online” RL settings where an agent interacts with the environment while learning its policy. A major advantage of safe “online” RL is that policies can be trained from scratch without data collected previously. On the other hand, “offline” RL (Levine et al., 2020) is a framework to train a policy from a fixed amount of pre-collected data, potentially solving the fundamental issues regarding safety or risk. Offline RL is well-suited in the context of safe RL because the agent does not interact with the environment during training; thus, the policy training does not essentially pose any risk. Hence, safe offline RL has been and will be considered a promising and fundamental solution for achieving safety in RL. Although most of the existing safe offline RL literature is based on Problem 1 (e.g., Le et al. (2019), Lee et al. (2021)) as shown in Table 1, such research direction is expected to expand to other problem settings.
6 Conclusion
This paper provides a survey of safe RL based on constrained criteria from the perspective of formulations. In Section 3, we first present seven common constraint formulations and associated representative algorithms. Also, the curated selection of literature is summarized according to the problem formulation in Table 1. In Section 4, we discuss the theoretical relations among formulations, which will be helpful for readers to obtain a deeper understanding of safe RL. Finally, we describe the current status of safe RL research and potential future directions. With this survey paper, we hope to have presented a systematic understanding and encourage further fundamental and applied research in this area.

### v2

@@
 Abstract
-Ensuring safety is critical when applying reinforcement learning (RL) to real-world problems.
-Consequently, safe RL emerges as a fundamental and powerful paradigm for safely optimizing an agent’s policy from experimental data.
-A popular safe RL approach is based on a constrained criterion, which solves the problem of maximizing expected cumulative reward under safety constraints.
-Though there has been recently a surge of such attempts to achieve safety in RL, a systematic understanding of the field is difficult due to 1) the diversity of constraint representations and 2) little discussion of their interrelations.
-To address this knowledge gap, we provide a comprehensive review of representative constraint formulations, along with a curated selection of algorithms specifically designed for each formulation.
-Furthermore, we elucidate the theoretical underpinnings that reveal the mathematical mutual relations among common problem formulations.
+Safety is critical when applying reinforcement learning (RL) to real-world problems.
+As a result, safe RL has emerged as a fundamental and powerful paradigm for optimizing an agent’s policy while incorporating notions of safety.
+A prevalent safe RL approach is based on a constrained criterion, which seeks to maximize the expected cumulative reward subject to specific safety constraints.
+Despite recent effort to enhance safety in RL, a systematic understanding of the field remains difficult.
+This challenge stems from the diversity of constraint representations and little exploration of their interrelations.
+To bridge this knowledge gap, we present a comprehensive review of representative constraint formulations, along with a curated selection of algorithms designed specifically for each formulation.
+In addition, we elucidate the theoretical underpinnings that reveal the mathematical mutual relations among common problem formulations.
 We conclude with a discussion of the current state and future directions of safe reinforcement learning research.
 1 Introduction
-Reinforcement learning (RL, Sutton and Barto (1998)) is a powerful paradigm for solving sequential decision-making problems through evaluation and improvement, which has made significant progress in many fields such as games (Mnih et al., 2015; Silver et al., 2016), robotics (Levine et al., 2016), and recommendation systems (Afsar et al., 2022).
+Reinforcement learning (RL, (Sutton and Barto, 1998)) is a powerful paradigm for solving sequential decision-making problems through evaluation and improvement, which has made significant progress in many fields such as games (Mnih et al., 2015; Silver et al., 2016; Wurman et al., 2022), robotics (Levine et al., 2016), data center cooling (Li et al., 2019), finance (Hambly et al., 2023), recommendation systems (Afsar et al., 2022), and healthcare (Yu et al., 2021).
 Recently, particular attention has been paid to RL for its use in the fine-tuning of large language models (LLMs) under the name of reinforcement learning from human feedback (RLHF, Ouyang et al.
 (2022)).
-RL is a general concept that can be used in many domains and has been spreading to various disciplines as a useful tool.
-When applying RL to real problems, safety is an essential requirement (Amodei et al., 2016).
-Promising areas for safe RL applications include robotics (Pham et al., 2018), autonomous driving (Shalev-Shwartz et al., 2016), and medical care (Jia et al., 2020), and so onL4D.
+RL is a general concept that can be employed in many domains and has been spreading to various disciplines as a useful paradigm.
+When applying RL to real-world problems, safety is an essential requirement (Amodei et al., 2016).
+Thus, safe RL has been actively studied in recent years so that the benefits of RL are realized while minimizing negative safety issues.
+Promising areas for safe RL applications include robotics (Pham et al., 2018), autonomous driving (Shalev-Shwartz et al., 2016), healthcare (Jia et al., 2020), and many others.
 An emerging and crucial use case of safe RL involves refining LLMs using RLHF to align with human preferences.
-It is critical to prevent harmful outputs (e.g., toxicity, discrimination) while maintaining utility, which is called Safe RLHF (Dai et al., 2023).
-As such, safe RL has been an active topic in artificial intelligence, which has been studied both theoretically and practically.
-Safe RL is a broad concept with different formulations for the different aspects of real-world safety-critical problems.
-Garcıa and Fernández (2015) provided a survey of safe RL and categorized its optimization criteria into four groups: 1) the constrained criterion (Geibel, 2006), 2) the worst-case criterion (Heger, 1994), 3) the risk-sensitive criterion (Borkar, 2002), and 4) others (e.g., r-squared and value-at-risk).
-This paper focuses on the constrained criterion of safe RL.
-It is a growing and powerful group of methods that maximize the expected cumulative reward under safety constraints.
-An interesting yet confusing point in the safe RL research based on the constrained criterion is that there are diverse representations of safety constraints yet little discussion regarding the relations among them.
+Specifically, it is critical to prevent harmfulness or bias (e.g., toxicity, discrimination) while maintaining the helpfulness of the generated sentences.
+For example, Safe RLHF (Dai et al., 2023) is a representative approach to balance such a tradeoff.
+Safe RL is an active area of research in artificial intelligence, with extensive theoretical and practical investigations aimed at developing RL systems that exhibit safe and reliable behavior.
+Safe RL is inherently a broad concept with different formulations for the different aspects of real-world safety-critical problems.
+Garcıa and Fernández (2015) provided an eminent survey of safe RL and categorized its optimization criteria into four groups: 1) constrained criteron (Geibel, 2006), 2) worst-case criteron (Heger, 1994), 3) risk-sensitive criteron (Borkar, 2002), and 4) others (e.g., r-squared, value-at-risk).
+This paper focuses on the constrained criterion of safe RL as a growing and powerful group of methods to optimize policies under safety constraints.
+A noteworthy aspect of safe reinforcement learning research based on constrained optimization criteria is the existence of diverse representations for safety constraints, with little analysis the interrelationships and theoretical connections among these various formulations.
 Safe RL research in the last decade has focused on developing new algorithms, i.e.
@@
 New algorithms have continuously been developed under various formulations, thereby making it increasingly challenging to stay abreast of advancements in the field.
-There are several recent survey papers on safe RL (Brunke et al., 2021; Liu et al., 2021; Gu et al., 2022), but they also focus on methods rather than formulations.
-As formulation represents the initial phase in comprehending safe RL or implementing algorithms in practical scenarios, it is imperative for the community to comprehensively survey existing literature and lay the groundwork for acquiring a systematic comprehension.
+There are several recent survey papers on safe RL (Kim et al., 2020; Brunke et al., 2021; Liu et al., 2021; Gu et al., 2022; Zhao et al., 2023b), but they also focus on methods rather than formulations.
+As formulation represents the initial phase in comprehending safe RL or implementing algorithms in practical scenarios, it becomes imperative for the community to comprehensively survey the existing literature and lay the groundwork for acquiring a systematic comprehension.
 Our contributions.
-This paper first provides a comprehensive survey focusing on the constraint formulations in safe reinforcement learning and introduces representative algorithms for each formulation.
-Also, we discuss the relations between each constraint formulation while defining three notions called transformability, generalizability, and conservative approximation.
-Specifically, we provide theoretical results that there are two problems called identical or more general safe RL (IoMG-SafeRL) problems, into which other common problems can be transformed.
-The main contribution of this paper is to bridge the problems faced by the reader with appropriate algorithms by organizing existing research with a focus on constraint formulation.
+This paper provides a comprehensive survey focusing on constraint formulations in safe reinforcement learning and introduces representative algorithms for each formulation.
+Furthermore, we discuss the relationships between various constraint formulations by defining three theoretical notions: transformability, generalizability, and conservative approximation.
+Specifically, we present theoretical results demonstrating that there exist two problems, termed Identical or More General Safe RL (IoMG-SafeRL) problems, into which other common problems can be either transformed or conservatively approximated.
+The main contribution of this paper is to bridge the gaps between the safe RL problems with appropriate algorithms by organizing existing research with a focus on constraint formulation.
 2 Preliminaries
@@
 $\mathcal{M}\cup\mathcal{C}\coloneqq\underbrace{\langle\,\mathcal{S},\mathcal{A},\mathcal{P},H,r,\gamma_{r},\rho\,\rangle}_{\text{Standard MDP}\,(\mathcal{M})}\cup\ \mathcal{C},$ (1)
-where $\mathcal{S}$ is a state space $\{s\}$, $\mathcal{A}$ is an action space $\{a\}$, and $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the state transition, where $\mathcal{P}(s^{\prime}\,|\,s,a)$ is the probability of transition from state $s$ to state $s^{\prime}$ when action $a$ is taken.
+where $\mathcal{S}\coloneqq\{s\}$ is a state space, $\mathcal{A}\coloneqq\{a\}$ is an action space, and $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the state transition, where $\mathcal{P}(s^{\prime}\,|\,s,a)$ is the probability of transition from state $s$ to state $s^{\prime}$ when action $a$ is taken.
 Note that $\Delta(X)$ denotes the probability simplex over the set $X$.
@@
 An interesting yet complicated point for understanding safe RL research is that there are several formulations for representing the constrained tuple $\mathcal{C}$ and the feasible policy space $\widehat{\Pi}$ depending on the problem settings or applications that researchers have in their minds.
-In the next section, we will review seven common safety constraint representations that have been well-studied in existing safe RL research.
+In the next section, we will review seven common safety constraint representations that have been well-studied in safe RL literature.
 […]
@@
 3.6 Almost Surely Instantaneous Safety Constraint with Time-invariant Threshold
-Some existing studies formulate safe RL problems via an instantaneous constraint, attempting to ensure safety even during the learning stage while aiming for extremely safety-critical applications such as planetary exploration (Turchetta et al., 2016; Wachi et al., 2018; Wachi and Sui, 2020; Wang et al., 2023).
+Some existing studies formulate safe RL problems via an instantaneous constraint, attempting to ensure safe exploration (Sui et al., 2015) even during the learning stage while aiming for extremely safety-critical RL applications such as planetary exploration and medical treatment (Turchetta et al., 2016; Wachi et al., 2018; Wachi and Sui, 2020; Wang et al., 2023).
 Such studies require the agent to satisfy the following instantaneous safety constraint at every time step.
@@
 $\max_{\pi}V^{\pi}_{r}(\rho)\quad\text{s.t.}\quad\textrm{Var}\left[\,V_{r}^{\pi}(\rho)\,\right]<\xi.$
-Conventional RL algorithms depend on ergodicity assumption; that is, any state $s$ is eventually reachable from any other state $s^{\prime}$ by following a suitable policy.
+The variance of the return is related to the Sharpe ratio (Sharpe, 1966) – the ratio between the expected profit and its standard deviation.
+Thus, this formulation is particularly useful in financial appliactions (Meng and Khushi, 2019).
+Also, conventional RL algorithms depend on ergodicity assumption; that is, any state $s$ is eventually reachable from any other state $s^{\prime}$ by following a suitable policy.
 This assumption does not hold in many practical applications since an agent cannot recover on its own after catastrophic actions.
@@
 While this paper has dealt with numerical safety constraints, there have been attempts to represent them using a certain language.
-For example, Fulton and Platzer (2018) represents the safety constraint via formal languages.
-This constraint representation is quite useful for leveraging human knowledge in RL, which leads to powerful solutions such as shielding methods (Alshiekh et al., 2018).
+For example, Fulton and Platzer (2018) or Hasanbeig et al.
+(2020) represent the safety constraint via formal languages such as linear temporal logic.
+This constraint representation is quite useful for leveraging human knowledge in RL, which leads to powerful solutions such as shielding methods (Alshiekh et al., 2018; Li and Bastani, 2020).
 Finally, Yang et al.
 (2021) uses natural language to represent safety constraints.
-This approach is friendly for the general public, which would be a promising method for applying safe RL to real problems.
+Natural language is one of the most powerful mediums yet friendly representations to the general public.
+Given the recent remarkable progress and utilities of LLMs, this approach would be promising for applying safe RL to real-world AI systems.
 Problem Type Representative Work Algorithm Theoretical Guarantee Open Source Software (OSS)
@@
 Amani et al. (2021) SLUCB Y T and C --
+Zhao et al. (2023a) SCPO Y C --
 Offline Amani and Yang (2022) Safe-DPVI Y T and C --
@@
 Problem 7 Online Shi et al. (2023) LSVI-NEW Y T and C --
-Wachi et al. (2023) MASE Y T and C A
+Wachi et al. (2023) MASE Y T and C --
 Table 1: Common safe RL formulations based on the constrained criterion and associated representative work.
@@
 Therefore, Problem 5 is identical to a special case of Problem 7 with $\xi_{h}\coloneqq\eta_{h}$ for all $h\in[H]$. Therefore, we obtain the desired lemma. ∎
-4.3 Two IoMG-SafeRL Problems
+4.3 The Two IoMG-SafeRL Problems
 We now provide three theorems on the interrelations among the common safe RL problems.
@@
 5.1 Formulation and Algorithm Selection
-As we have discussed in this paper, when we adopt the safe RL paradigm, there are two main steps: 1) problem formulation and 2) policy optimization.
+As mentioned earlier, when we adopt the safe RL paradigm, there are two main steps: 1) problem formulation and 2) policy optimization.
 Based on the survey results, let us discuss how to solve safe RL problems via constraints.
@@
 Another perspective in the algorithm selection should include when an agent needs to satisfy the safety constraint.
-Existing algorithms can be divided into two classes (see “safety” column in Table 1).
+Existing algorithms can be divided into two classes (see “safety” column in Table 1.
 The first class (w/o T) tries to achieve the required level of safety after convergence while encouraging safety during training.
@@
 On the other hand, “offline” RL (Levine et al., 2020) is a framework to train a policy from a fixed amount of pre-collected data, potentially solving the fundamental issues regarding safety or risk.
-Offline RL is well-suited in the context of safe RL because the agent does not interact with the environment during training; thus, the policy training does not essentially pose any risk.
-Hence, safe offline RL has been and will be considered a promising and fundamental solution for achieving safety in RL.
+Offline RL is well-suited in the context of safe RL because the agent does not interact with the real environment during training; thus, the policy training does not essentially pose any risk.
+Hence, safe offline RL is a promising approach to achieve safety in RL if enough data or high-fidelity simulation is available.
 Although most of the existing safe offline RL literature is based on Problem 1 (e.g., Le et al.
@@
 6 Conclusion
-This paper provides a survey of safe RL based on constrained criteria from the perspective of formulations.
-In Section 3, we first present seven common constraint formulations and associated representative algorithms.
-Also, the curated selection of literature is summarized according to the problem formulation in Table 1.
-In Section 4, we discuss the theoretical relations among formulations, which will be helpful for readers to obtain a deeper understanding of safe RL.
-Finally, we describe the current status of safe RL research and potential future directions.
-With this survey paper, we hope to have presented a systematic understanding and encourage further fundamental and applied research in this area.
+This paper provides a comprehensive survey of safe reinforcement learning based on constrained optimization criteria, with a particular focus on problem formulations.
+We present seven common constraint formulations and their associated representative algorithms in Section 3.
+Additionally, a curated selection of relevant literature is summarized according to the problem formulation in Table 1.
+In Section 4, we examine the theoretical relations among these formulations, offering readers a deeper understanding of safe RL.
+Furthermore, we describe the current status of safe RL research and outline potential future directions.
+Through this survey paper, we aim to foster a systematic understanding of constraint formulations and encourage further fundamental and applied research in safe reinforcement learning.